6n + 4 = 16, find the value of 5n - 7


1
Expert's answer
2021-05-07T11:56:30-0400

"Solution: Given ~that :6n+4=16
\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~6n=16-4
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~6n=12
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~n=\frac{12}{6}
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~n=2
\\~~~~~~~~~~~~~~~~~~~\therefore5n-7=5(2)-7 
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=10-7
\\~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=3
\\~~~~~~~~~~~~~~~~~~~~\therefore 5n-7=3"